# Natural Counterfactuals With Necessary Backtracking

### v2

@@ Abstract
Counterfactual reasoning is pivotal in human cognition and especially important for providing explanations and making decisions. While Judea Pearl’s influential approach is theoretically elegant, its generation of a counterfactual scenario often requires interventions that are too detached from the real scenarios to be feasible. In response, we propose a framework of natural counterfactuals and a method for generating counterfactuals that are natural with respect to the actual world’s data distribution. Our methodology refines counterfactual reasoning, allowing changes in causally preceding variables to minimize deviations from realistic scenarios. To generate natural counterfactuals, we introduce an innovative optimization framework that permits but controls the extent of backtracking with a “naturalness” criterion. Empirical experiments indicate the effectiveness of our method.
@@ 1 Introduction
Counterfactual reasoning, which aims to answer what a feature of the world would have been if some other features had been different, is often used in human cognition, to perform self-reflection, provide explanations, and inform decisions. For AI systems to mirror such human-like decision-making processes, incorporating counterfactual reasoning is crucial. Judea Pearl’s structural approach to counterfactual modeling and reasoning stands as a cornerstone in machine learning (Pearl, 2009). Within this framework, counterfactuals are conceptualized as being generated by surgical interventions on the features to be changed that sever the original causal links responsible for those features, while leaving its causally upstream features untouched. Such non-backtracking counterfactual reasoning (i.e., reasoning about the consequences of an change without tracing back to changes in causally preceeding variables) can yield valuable insights into the consequences of hypothetical actions. Consider a scenario: a sudden brake of a high-speed bus caused Tom to fall and injure Jerry, as illustrated in Fig. 1. Non-backtracking counterfactual reasoning would tell us that if Tom had stood still (despite the sudden braking), then Jerry would not have been injured. Pearl’s approach supplies a principled machinery to reason about conditionals of this sort, which are usually useful for explanation, planning, and responsibility allocation.
Figure 1: Motivational Example: Treating Bus, Tom, and Jerry as Variables. Downward arrows indicate their values, while upward arrows represent interventions.
@@
However, such surgical interventions may not be feasible in practice and may not provide significant assistance in self-reflection. In the previous example, preventing Tom’s fall in a sudden braking scenario requires defying mechanisms that are difficult or physically impossible to disrupt, such as the law of inertia. This supposed hard intervention may thus be too far-fetched to be relevant for practical purposes. For example, from a legal perspective, Tom’s fall causing Jerry’s injury could be given a “necessity defense,” acknowledging that the sudden braking left him with no alternatives (Conde, 1981). Hence, for the purpose of allocating responsibility, reasoning about the counterfactual situation of Tom standing still despite the sudden braking is arguably irrelevant or even misleading.
Our paper addresses counterfactual reasoning with a focus on outcomes that are constructive, which are intended to enhance practical situations and offer actionable insights. To achieve this, we ensure that every change in a system is natural. Thus, after these modifications, the new counterfactual data point will be natural with respect to data distribution in the actual world. Accordingly, we introduce the notion of “natural counterfactuals” to address the limitations of non-backtracking counterfactuals discussed above. For example, as depicted in Fig. 1, rather than the infeasible scenario where Tom does not fall at a sudden bus stop, a more realistic intervention would involve the bus slowing down earlier, achieved through backtracking (Lewis, 1979) that keeps the scenario within real-world bounds. Furthermore, one must ensure that the counterfactual data point remains as close as possible to the original data point; otherwise, unnecessary interventions may affect too many variables. To address this concern, we formulate a minimal change principle that guides us in performing only the necessary backtracking.
@@
Moreover, from a machine learning perspective, when hard interventions lead to unrealistic scenarios relative to the training data, predicting counterfactual outcomes in such scenarios can be highly uncertain and inaccurate (Hassanpour & Greiner, 2019). This issue becomes particularly pronounced when non-parametric models are employed, as they often struggle to generalize to unseen, out-of-distribution data (Schölkopf et al., 2021). The risk of relying on such counterfactuals is thus substantial, especially in high-stake applications like healthcare and autonomous driving. To tackle this, we develop an approach that amounts to utilizing only feasible interventions that keep data within its original distribution, allowing backtracking when needed. This strategy effectively reduces the risk of inaccurate predictions and ensures more reliable results.
@@
Consequently, our approach aims to achieve the goal of ensuring that counterfactual scenarios remain sufficiently realistic with respect to the actual world’s data distribution by permitting minimal yet necessary backtracking. It is designed on two major considerations: First, we need criteria to determine the feasibility of interventions, ensuring they are realistic with respect to actual-world data distribution. Second, we appeal to backtracking when and only when it is necessary to avoid infeasible interventions, and need to develop a feasible optimization framework to realize this strategy. The key contributions of this paper include:
@@ 2 Related Work
Non-backtracking Counterfactual Generation. As will become clear, our theory is presented in the form of counterfactual sampling or generation. (Ribeiro et al., 2023; Kocaoglu et al., ; Dash et al., 2022; Sanchez & Tsaftaris, ) use the deep generative models to learn an SCM from data given a causal graph; these works strictly follow Pearl’s theory of non-backtracking counterfactuals. Our case studies will examine some of these models and demonstrate their difficulties in dealing with interventions that are unrealistic relative to training data.
Backtracking Counterfactuals. Backtracking in counterfactual reasoning has drawn plenty of attention in philosophy (Hiddleston, 2005), psychology (Dehghani et al., 2012), and cognitive science (Gerstenberg et al., 2013). Hiddleston (2005) proposes a theory that is in spirit similar to ours, in which backtracking is allowed but limited by some requirement of matching as much causal upstream as possible. Gerstenberg et al. (2013) shows that people use both backtracking and non-backtracking counterfactuals in practice and tend to use backtracking counterfactuals when explicitly required to explain causes for the supposed change in a counterfactual. von Kügelgen et al. (2022) is a most recent paper explicitly on backtracking counterfactuals. The main differences between that work and ours are that von Kügelgen et al. (2022) requires backtracking all the way back to exogenous noises and measures closeness on noise terms, which are less desirable than limiting backtracking to what we call “necessary backtracking” and measuring closeness directly on endogenous, observable variables, because changes to the unobserved or noise terms are by definition outside of our control and not actionable. Moreover, their backtracking counterfactuals sometimes allow gratuitous changes, as we explain in Sec. G of the Appendix.
Counterfactual Explanations. A prominent approach in explainable AI is counterfactual explanation (Wachter et al., 2018; Dhurandhar et al., 2018; Mothilal et al., 2020; Barocas et al., 2020; Verma et al., 2020; Schut et al., 2021), on which our work is likely to have interesting bearings. Most works on this topic define some sense of minimal changes of an input sample with a predicted class $c$ such that adding the minimal changes into the input would make it be classified into another (more desirable) class. Although this paper does not discuss counterfactual explanations, our framework may well be used to define a novel notion of counterfactual explanation by requiring the counterfactual instances to be “natural” in our sense.
3 Notations and Background
@@
Structural Causal Models. We use a structural causal model (SCM) to represent the data generating process of a causal system. A SCM is a mathematical structure consisting of a triplet ${\mathcal{M}}:=<{\bf U},{\bf V},{\bf F}>$, with an exogenous (noise) variable set ${\bf U}=\{{\bf U}_{1},...,{\bf U}_{N}\}$, an endogenous (observed) variable ${\bf V}=\{{\bf V}_{1},...,{\bf V}_{N}\}$, and a function set ${\bf f}=\{f_{1},...,f_{N}\}$ (Pearl, 2009). Each function, $f_{i}\in{\bf f}$, specifies how an endogenous variable ${\bf V}_{i}$ is determined by its parents ${\mathbf{PA}}_{i}\subseteq{\bf V}$:
@@
Local Mechanisms. A local mechanism is the conditional distribution of an endogenous variable based on its parent variables, represented as $p({\bf V}_{i}|{\mathbf{PA}}_{i})$ for $i=1,...,N$. This definition inherently captures the properties of noise variables; given a fixed value for ${\mathbf{PA}}_{i}$, noise ${\bf U}_{i}$ entirely dictates the probability of ${\bf V}_{i}$ (Pearl, 2009). Hence, throughout this paper, the term “local mechanisms” will encompass both the conditional distribution of an endogenous variable given its parents $p({\bf V}_{i}|{\mathbf{PA}}_{i})$ and the distribution of the noise variable $p({\bf U}_{i})$.
Intervention. Given an SCM, an intervention on an endogenous variable ${\bf A}$ is represented by replacing its function with a constant function ${\bf A}={\bf a}^{*}$, where ${\bf a}^{*}$ is the target value of ${\bf A}$, and leaving functions and local mechanisms for other variables intact (Pearl, 2009).
Non-Backtracking Counterfactuals. A counterfactual ponders what would happen in a scenario that differs from the actual one in a certain way. Following a standard notation, terms with a *** superscript refer to a counterfactual world. For example, ${\bf u}^{*}_{i}$ denotes ${\bf U}_{i}$’s value in a counterfactual world. Let ${\bf A}$, ${\bf B}$, and ${\bf E}$ be endogenous variables. Here is a general counterfactual question: given evidence ${\bf E}={\bf e}$, what would the value of ${\bf B}$ have been if ${\bf A}$ had been ${\bf a}^{*}$? The Pearlian, non-backtracking reading of this question takes the counterfactual supposition of ${\bf A}={\bf a}^{*}$ to be realized by an intervention on ${\bf A}$ (Pearl, 2009). Given this understanding, counterfactual inference includes three steps. (1) Abduction: The noise distribution is updated based on the evidence ${\bf E}={\bf e}$. (2) Action: The causal model is modified, in which ${\bf A}$ is fixed to ${\bf a}^{*}$ while keeping other components the same as before. (3) Prediction: The counterfactual outcome of ${\bf B}$ is inferred using the updated noise distribution and the modified model.
@@ 4 A Framework for Natural Counterfactuals
In considering a counterfactual scenario in which ${\bf A}$’s value is ${\bf a}^{*}$ instead of ${\bf a}$, the non-backtracking mode always appeals to a direct intervention on ${\bf A}$, i.e., $do({\bf A}={\bf a}^{*})$ in Pearl’s influential notation. However, in our framework, the counterfactual supposition is not necessarily realized by a direct intervention on ${\bf A}$, and may be realized by interventions on some causally preceding variables, depending on the “naturalness” of interventions. To differentiate from the intervention $do({\bf A}={\bf a}^{*})$ in non-backtracking counterfactuals, we introduce a specific operator $change(\cdot)$ and $change({\bf A}={\bf a}^{*})$ signifies a desired modification in ${\bf A}$, which may result from an intervention on ${\bf A}$ or an intervention on some causal ancestors of ${\bf A}$.
We introduce two new concepts to determine where to carry out feasible interventions on ${\bf C}$ (either ${\bf A}$ itself or the causal earlier variables of ${\bf A}$) to realize $change({\bf A}={\bf a}^{*})$ in natural counterfactuals: a naturalness constraint and a necessary backtracking principle. [^1] Given a feasible intervention, our natural counterfactual inference also follows three steps: abduction, action, and prediction, aligned with the procedure of non-backtracking counterfactuals as mentioned in Sec. 3. More details for the procedure of natural counterfactuals are in Sec. E. First, we propose using a naturalness criterion to assess the feasibility of an intervention. We assume that actual-world data contain all relevant causal mechanisms for a given causal graph; thus, out-of-distribution data are physically infeasible. This means we limit changed values resulting from interventions to stay within the observed data distribution, ensuring the feasibility of interventions on ${\bf C}$ and the naturalness of resulting counterfactuals.
Furthermore, we consider the proximity between the actual value and its counterfactual value, leading to a principle known as “necessary backtracking.” This principle aims to minimize changes in the counterfactual world and backtrack as little as possible, keeping ${\bf C}$ as closely aligned with ${\bf A}$ as feasible. In summary, the naturalness constraint and necessary backbreaking both contribute to determining a feasible intervention.
4.1 Feasible Intervention Optimization
To address the challenge of determining a feasible intervention, we propose treating it as an optimization problem. This entails optimizing a defined distance metric between the counterfactual outcome produced by the intervention and the corresponding real-world data point, following the principle of necessary backtracking. This optimization procedure also enforces a naturalness constraint on counterfactuals. We define this optimization framework as Feasible Intervention Optimization (FIO) as follows:
@@ 4.1 Feasible Intervention Optimization
where $an({\bf A})$ and $an({\bf A})^{*}$ represent the actual value and counterfactual value of ${\bf A}$’s ancestors ${\mathbf{AN}}({\bf A})$ respectively, where ${\bf A}\in{\mathbf{AN}}({\bf A})$. $D(\cdot)$ is a specific distance metric to assess the distance between actual world and counterfactual world. $g_{n}(\cdot)$ measures the naturalness of counterfactual value of ${\bf A}$’s ancestors and $\epsilon$ is a small constant. We will discuss $D(\cdot)$ and $g_{n}(\cdot)$ later.
Optimization Scope. All possible backtracking variables are contained within ${\bf A}$’s ancestor set ${\mathbf{AN}}({\bf A})$. Consequently, our optimization focus remains confined to these nodes. Through FIO, we obtain the optimal value $an({\bf A})^{*}$. This identifies the changed variables, ${\bf C}$, which we target for intervention, enabling the definition of the feasible intervention $do({\bf C}={\bf c}^{*})$, with ${\bf c}^{*}$ as the post-optimization counterfactual values of ${\bf C}$.
Naturalness Constraint. In addition to ensuring the change ${\bf A}={\bf a}^{*}$, the counterfactual value $an({\bf A})^{*}$ must satisfy a naturalness criterion to ensure the feasibility of interventions. Specifically, its naturalness should exceed the threshold $\epsilon$ to remain within the data distribution. A detailed discussion of potential naturalness constraints is presented in Sec. 4.2.
Distance Measure $D(\cdot)$. $D(\cdot)$ considers the concept of necessary backtracking and serves as a metric for quantifying the distance between the counterfactual and actual worlds. For a detailed exploration, please refer to Sec. 4.3.

### v3

@@ Abstract
Counterfactual reasoning is pivotal in human cognition and especially important for providing explanations and making decisions. While Judea Pearl’s influential approach is theoretically elegant, its generation of a counterfactual scenario often requires too much deviation from the observed scenarios to be feasible, as we show using simple examples. To mitigate this difficulty, we propose a framework of natural counterfactuals and a method for generating counterfactuals that are more feasible with respect to the actual data distribution. Our methodology incorporates a certain amount of backtracking when needed, allowing changes in causally preceding variables to minimize deviations from realistic scenarios. Specifically, we introduce a novel optimization framework that permits but also controls the extent of backtracking with a “naturalness” criterion. Empirical experiments demonstrate the effectiveness of our method. The code is available at https://github.com/GuangyuanHao/natural_counterfactuals.
@@ 1 Introduction
Counterfactual reasoning, which aims to answer what a feature of the world would have been if some other features had been different, is often used in human cognition, to perform self-reflection, provide explanations, and inform decisions (Roese, 1997; Byrne, 2016). For AI systems to achieve human-like abilities of reflection and decision-making, incorporating counterfactual reasoning is crucial. Judea Pearl’s structural approach to counterfactual modeling and reasoning (Pearl, 2009) has been especially influential in recent decades. Within this framework, counterfactuals are conceptualized as being generated by surgical interventions on the variables to be changed, while leaving its causally upstream variables intact and all downstream causal mechanisms invariant. These counterfactuals are thoroughly non-backtracking in that the desired change is supposed to happen without tracing back to changes in causally preceding variables. Reasoning about these counterfactuals can yield valuable insights into the consequences of hypothetical actions. Consider a scenario: a sudden brake of a high-speed bus caused Tom to fall and injure Jerry. Non-backtracking counterfactuals would tell us that if Tom had stood still (despite the sudden braking), then Jerry would not have been injured. Pearl’s approach supplies a principled machinery to reason about conditionals of this sort, which are usually useful for explanation, planning, and responsibility allocation.
@@
However, such surgical interventions are sometimes so removed from what are or can be observed that it is difficult or even impossible to learn from data the consequences of such interventions. In the previous example, preventing Tom’s fall in a sudden braking scenario requires defying mechanisms that are difficult or even physically impossible to disrupt. As a result, there are likely to be no data points in the reservoir of observed scenarios that are consistent with a person standing still during a sudden braking. If so, it can be very challenging to learn to generate such a counterfactual from the available data, as we will demonstrate in our experiments. [^1] Moreover, one may argue that the envisaged counterfactual may be too unnatural to be relevant for practical purposes. For example, from a legal perspective, Tom’s causing Jerry’s injury could be given a “necessity defense,” acknowledging that the sudden braking left him with no alternatives (Conde, 1981). Hence, for the purpose of allocating responsibility, reasoning about the counterfactual situation of Tom standing still despite the sudden braking is perhaps irrelevant or even misleading.
In this paper, we introduce a notion of “natural counterfactuals” to address the above issue with non-backtracking counterfactuals. Our notion will allow a certain amount of backtracking, to keep the counterfactual scenario “natural” with respect to the available observations. For example, rather than the unrealistic scenario where Tom does not fall at a sudden bus stop, a more natural counterfactual scenario to realize the change to not-falling would involve changing at the same time some causally preceding events, such as changing the sudden braking to gradually slowing down. On the other hand, our notion also constrains the extent of backtracking; in addition to a naturalness criterion, we formulate an optimization scheme to encourage minimizing backtracking while meeting the naturalness criterion.
@@
As we will show empirically, this new notion of counterfactual is especially useful from a machine learning perspective. When interventions lead to unrealistic scenarios relative to the training data, predicting counterfactual outcomes in such scenarios can be highly uncertain and inaccurate (Hassanpour & Greiner, 2019). This issue becomes particularly pronounced when non-parametric models are employed, as they often struggle to generalize to unseen, out-of-distribution data (Schölkopf et al., 2021). The risk of relying on such counterfactuals is thus substantial, especially in high-stake applications like healthcare and autonomous driving. In contrast, our approach amounts to searching for feasible changes that keep generated counterfactuals within its original distribution, employing backtracking when needed. This strategy effectively reduces the risk of inaccurate predictions and ensures more reliable results.
@@
In short, our approach aims to achieve the goal of ensuring that counterfactual scenarios remain sufficiently realistic with respect to the actual data distribution by permitting minimal yet necessary backtracking. It is designed with two major elements. First, we need criteria to determine the feasibility of interventions, ensuring they are realistic with respect to the actual data distribution. Second, we appeal to backtracking when and only when it is necessary to avoid infeasible interventions, and need to develop an optimization framework to realize this strategy. To be clear, our aim is not to propose a uniquely correct semantics for counterfactuals or to show that our notion of counterfactuals is superior to others for all purposes. Rather, our goal is to develop a notion of counterfactual that is theoretically well-motivated and practically useful in some contexts and for certain purposes. There may in the end be a general semantics to unify fruitfully our notion and other notions of counterfactual in the literature, but we will not attempt that in this paper. Our working assumption is that there can be different but equally coherent notions of counterfactuals, and we aim to show that the notion developed in this paper is particularly useful for generating counterfactual instances that are feasible with respect to data.
The key contributions of this paper include:
@@ 2 Related Work
Non-backtracking Counterfactual Generation. As will become clear, our theory is presented in the form of counterfactual sampling or generation. (Ribeiro et al., 2023; Kocaoglu et al., ; Dash et al., 2022; Sanchez & Tsaftaris, ) use the deep generative models to learn a causal model from data given a causal graph, and then use the model to generate non-backtracking counterfactuals. Our experiments will examine some of these models and demonstrate their difficulties in dealing with interventions that are unrealistic relative to training data, due to the fact that the causal model learned from data is not reliable in handling inputs that are out-of-distribution.
Backtracking Counterfactuals. Backtracking in counterfactual reasoning has drawn plenty of attention in philosophy (Hiddleston, 2005), psychology (Dehghani et al., 2012), and cognitive science (Gerstenberg et al., 2013). (Hiddleston, 2005) proposes a theory that is in spirit similar to ours, in which backtracking is allowed but limited by some requirement of matching as much causal upstream as possible. (Gerstenberg et al., 2013) shows that people use both backtracking and non-backtracking counterfactuals in practice and tend to use backtracking counterfactuals when explicitly required to explain causes for the supposed change in a counterfactual. (von Kügelgen et al., 2022) is the first work, as far as we know, to formally introduce backtracking counterfactuals in the framework of structural causal models. The main differences between this work and ours are that (von Kügelgen et al., 2022) requires “full” backtracking given a causal model, all the way back to exogenous noise terms, and measures closeness in terms of these noise terms, whereas we limit backtracking to what is needed to render the counterfactual “natural”, and the measure of closeness in our framework can be defined directly on endogenous, observed variables, which is arguably desirable because changes to the unobserved variables are by definition outside of our control and not actionable. More detailed comparisions can be found in Sec. F of the Appendix.
Algorithmic Recourse and Counterfactual Explanations. Algorithmic recourse (AR) (Karimi et al., 2021) and counterfactual explanations (CE) (Wachter et al., 2018; Dhurandhar et al., 2018; Mothilal et al., 2020; Barocas et al., 2020; Pawlowski et al., 2020a; Verma et al., 2020; Schut et al., 2021; Yan & Wang, 2023) represent two leading strategies within explainable AI that heavily exploit counterfactual reasoning or reasoning about intervention effects. Various studies (Karimi et al., 2021; Artelt & Hammer, 2020; Poyiadzi et al., 2020; Joshi et al., 2019; Yan & Wang, 2023) explore concepts of feasibility akin to that of “naturalness” in this work, though apply them in quite different contexts. The objectives of CE and AR are to pinpoint minimal alterations to an input (in CE) or minimal interventions (in AR) that would either induce a desirable output from a predictive model (in CE) or lead to an desirable outcome (in AR). In contrast, our work here focuses primarily on generating counterfactuals. While this paper does not directly address CE or AR, our notion of natural counterfactuals is likely to be very relevant to these tasks.
3 Preliminaries
@@
Structural Causal Models (SCM). We assume there is an underlying recursive SCM (Pearl, 2009) of the following sort to represent the data generating process. A SCM ${\mathcal{M}}:=\langle{\bf U},{\bf V},{\bf f},p({\bf U})\rangle$ consists of exogenous (noise) variables ${\bf U}=\{{\bf U}_{1},...,{\bf U}_{N}\}$, endogenous (observed) variables ${\bf V}=\{{\bf V}_{1},...,{\bf V}_{N}\}$, functions ${\bf f}=\{f_{1},...,f_{N}\}$, and a joint distribution $p({\bf U})$ of noise variables, which are assumed to be jointly independent. Each function, $f_{i}\in{\bf f}$, specifies how an endogenous variable ${\bf V}_{i}$ is determined by its parents ${\mathbf{PA}}_{i}\subseteq{\bf V}$:
@@
Such a SCM entails a (causal) Bayesian network over the observed variables, consisting of the directed acyclic graph over ${\bf V}$ in which there is an arrow from each member of ${\mathbf{PA}}_{i}$ to ${\bf V}_{i}$, and the joint distribution of ${\bf V}$ induced by ${\bf f}$ and $p({\bf U})$. In our setting, we assume this causal graph is known and samples from this joint distribution are available, but ${\bf f}$ and $p({\bf U})$ are not given, though some assumptions on ${\bf f}$ and $p({\bf U})$ will be needed for identifiability.
Local Mechanisms. Functions in ${\bf f}$ are usually regarded as representing local mechanisms. In this paper, however, we will use the term “local mechanism” to refer to the conditional distribution of an endogenous variable given its parent variables, i.e., $p({\bf V}_{i}|{\mathbf{PA}}_{i})$ for $i=1,...,N$, which can be estimated from the available information. Note that a local mechanism in this sense implicitly encodes the properties of the corresponding noise variable; given a fixed value of ${\mathbf{PA}}_{i}$, the noise ${\bf U}_{i}$ determines the probability distribution of ${\bf V}_{i}$ (Pearl, 2009). Hence, the term “local mechanism” will also be used sometimes to refer to the distribution of the noise variable $p({\bf U}_{i})$.
Intervention. Given a SCM, an intervention on a set of endogenous variables ${\bf A}\subseteq{\bf V}$ is represented by replacing the functions for members of ${\bf A}$ with constant functions $X=x^{*}$, where $X\in{\bf A}$ and $x^{*}$ is the target value of $X$, and leaving the functions for other variables intact (Pearl, 2009). [^2] Following a standard notation, we use a * superscript to signal counterfactual values.
Non-Backtracking Counterfactuals. Let ${\bf A}$, ${\bf B}$, and ${\bf E}$ be sets of endogenous variables. A general counterfactual question takes the following form: given evidence ${\bf E}={\bf e}$, what would the value of ${\bf B}$ have been if ${\bf A}$ had taken the value setting ${\bf a}^{*}$? In this paper, we focus on a special case of this question in which ${\bf E}={\bf V}$, i.e., the evidence is a complete data point covering all observed variables. That is, given an actual date point, we consider what the data point would have been if a variable had taken a different value than its actual value. The Pearlian, non-backtracking reading of this question takes the counterfactual supposition of ${\bf A}={\bf a}^{*}$ to be realized by an intervention on ${\bf A}$ (Pearl, 2009). This means that in the envisaged counterfactual scenario, all variables in the causal upstream of ${\bf A}$ keep their actual values while ${\bf A}$ takes a different value. As mentioned, a potential problem is that such a scenario is outside of the support of available data, and so it is often unreliable to make inferences about the downstream variables in the scenario based on the available data.
@@ 4 A Framework for Natural Counterfactuals
Using Pearl’s influential do-operator, the non-backtracking mode appeals to $do({\bf A}={\bf a}^{*})$, an intervention to set the value ${\bf A}$ to ${\bf a}^{*}$, to generate a counterfactual instance. However, in our framework, the counterfactual supposition is not necessarily realized by an intervention on ${\bf A}$, while keeping all its causal upstream intact. Instead, when $do({\bf A}={\bf a}^{*})$ results in a counterfactual setting that violates a naturalness criterion, some backtracking will be invoked. To differentiate from the intervention $do({\bf A}={\bf a}^{*})$, we will often use $change(\cdot)$ and write $change({\bf A}={\bf a}^{*})$ to denote a desired modification in ${\bf A}$. Different semantics for counterfactuals correspond to different interpretations of the change-operator. In this paper, we explore an interpretation that connects the change-operator to the do-operator in a relatively straightforward manner.
The basic idea is that $change({\bf A}={\bf a}^{*})$ will correspond to $do({\bf C}={\bf c}^{*})$ for some set ${\bf C}$ that includes ${\bf A}$ and possibly some of ${\bf A}$’s causal ancestors. When ${\bf C}=\{{\bf A}\}$, this is equivalent to a non-backtracking interpretation. In general, however, some variables in ${\bf A}$’s causal upstream need to change together with ${\bf A}$ in order to keep the counterfactual scenario within the relevant support. A central component of our approach is to design a way to determine ${\bf C}$ and ${\bf c}^{*}$, given a request of $change({\bf A}={\bf a}^{*})$. We will call the resulting $do({\bf C}={\bf c}^{*})$ the least-backtracking feasible (LBF) intervention for $change({\bf A}={\bf a}^{*})$. Once the LBF intervention is determined, inferences can be made in the same fashion as in Pearl’s approach (Pearl, 2009).
To determine the LBF intervention for $change({\bf A}={\bf a}^{*})$, we formulate it as an optimization problem to search for a minimal change of ${\bf A}$’s causal ancestors that, together with changing ${\bf A}$ to ${\bf a}^{*}$, satisfy a “naturalness” criterion. Let ${\mathbf{AN}}({\bf A})$ denote the set of ${\bf A}$’s ancestors in the given causal graph together with ${\bf A}$ itself. We define the optimization framework, Feasible Intervention Optimization (FIO), as follows:
@@
where $an({\bf A})$ and $an({\bf A})^{*}$ represent the actual value setting and the counterfactual value setting of ${\mathbf{AN}}({\bf A})$ respectively (note that ${\bf A}\subseteq{\mathbf{AN}}({\bf A})$). $g_{n}(\cdot)$ measures the naturalness of the counterfactual value setting of ${\mathbf{AN}}({\bf A})$ and $\epsilon$ is a small constant. So the optimization has a naturalness criterion as a constraint. On the other hand, $D(\cdot)$ is a distance metric designed to encourage the counterfactual value setting to invoke the least amount of backtracking. Below we develop these two components in Sec. 4.1 and Sec. 4.2, respectively. Once we obtain the counterfactual value $an({\bf A})^{*}$ by FIO, the variable set for the LBF intervention includes ${\bf A}$ and other variables corresponding to the difference between $an({\bf A})^{*}$ and $an({\bf A})$, i.e., ${\bf A}$ is always included in ${\bf C}$ even when $an({\bf A})^{*}=an({\bf A})$.